# Persona-driven Simulation of Voting Behavior in the European Parliament with Large Language Models

### v2

@@ Appendix C Counterfactual Debates
Appendix D Proposals and Committees
We show the full list names of the proposals and the committees that initiated them in Table 5. We selected these proposals because they had a significant societal impact, evidenced by an official EU press release focused on that proposal decision and by the fact that they were the sole focus of one or more debates in the European Parliament. We analyzed the group line for each proposal and compared it with the predicted group line of our best performing approach Llama3-70B with reasoning and attribute prompts. We show the comparison in Table 6. We see that the model never predicts the ABSTENTION group line. The Renew group was predicted best, followed by S&D, Greens/EFA and EPP. For the far-left group GUE/NGL the model predicts more FOR group lines on topics that include Ukraine and policies that introduce more digital processes, which include more data collecting of citizens (e.g. European Health Data Space or Instant Payments in Euro). For the ECR group, the model predicts more agreement to proposals that try to preserve nature and reduce emissions (e.g. Nature Restoration, Methane emissions reduction in the energy sector). The ID group has the group line of ABSTENTION most often, which the model did not predict at all. Further false predictions are again related to ecological proposals, however for this group the model predicts more AGAINST group lines, than FOR group lines (e.g. Ozone Depleting Substances, Empowering consumers for the green transition). The vote breakdown indicates that our current models struggle to predict when politicians from centrist parties vote against their group’s established position. Conversely, for groups on the political fringes, the models sometimes fail to predict the group line itself.
@@ Appendix D Proposals and Committees
We analyzed the change of voting behavior across different policy categories when using counterfactual speeches as well. To explore this, we categorize proposals based on the committee that initiated them. Figure 6 reveals that individuals from left (S&D, Greens/EFA) and even far-left groups (GUE/NGL) consistently maintain their support for proposals on certain topics, such as those from the Employment and Social Affairs Committee and the Women’s Rights and Gender Equality Committee. For these proposals, most original speeches supported them ($83\%$ and $72\%$, respectively), so the majority of counterfactual debates advise against them. Despite this, the votes of individuals affiliated with left-wing groups remain unchanged. In contrast, center-right (Renew, EPP) and far-right groups (ECR, ID) are more likely to alter their votes on proposals from these committees.
@@
Appendix E Influence of different attributes
@@
E.1 Attribute Matching
We make use of regular expressions to determine which attributes appear in the reasoning chains of each MEP and for removing specific words from speeches. In the following, we list the substrings used to search for the values of each attribute in the reasoning chains. For each MEP, we only search for their actual attribute values (i.e. we search for Ireland and Irish if a MEP represents Ireland in the EU parliament). For the attributes age, we computed the age of each MEP at the time of voting. We used the exact wording for birthplace which comes from Wikidata [^6] https://www.wikidata.org/. And the exact wording for the politician’s name and national party from the European Parliament Website [^7] https://www.europarl.europa.eu/meps/en/full-list/all.
@@ E.1 Attribute Matching
E.2 Additional Attribute Analysis
@@
Table 8 shows the gender in the reasonings. We disaggregate the mentions of gender identifiers by gender of the MEPs (e.g. gender_fm indicates the percentage of reasonings by female (f) simulated MEPs that mention a male (m) identity word). We find that the percentages stay roughly consistent across all prompting strategies and find that female identity words are mentioned much more often than male identity words. Further, we observe that the reasoning chains contain male and female identity words, respectively, at a similar rate regardless of the gender in the prompt. We conclude that in the reasoning chains, gender identity terms are likely not used to refer to the gender of the persona (the simulated "own" gender), but rather to discuss a proposal related to gender (e.g. "Combating violence against women and domestic violence") or other proposals initiated by the Committee of Women’s Rights and Gender Equality.
Figure 7: Prompting example. This shows the different approaches we take to prompt the model. We give a persona description, which is either a Wikipedia Summary or an attribute prompt, proposal information in the form of debates (only real speeches are depicted here) and instructions on how to format the response (only no reasoning and reasoning are displayed here).
@@
Table 9 presents the number of politicians for whom the correct group is mentioned at least once in the reasoning when only their name is provided as persona. The pre-training data appears to contain information on a significant portion of politicians, though the distribution is biased toward center parties. Although the pre-training data likely contains more information than what is provided in the prompt, we still observe a significant change in behavior when including all attributes in the prompt instead of just the name. This suggests that using only the name is not always sufficient to guide the model toward a known public figure.
Appendix F Other models
@@
To prevent knowledge leakage we want to restrict used models to pretraining data cutoff before the date of our earliest vote in our dataset (16. January 2024) to the best of our abilitites and have comparable performance to LLama-3 and Qwen2.5. We identify two other models that fit this criteria: Phi-mini (Phi-3.5-mini-instruct abdin2024phi and (Mixtral-8x7B-Instruct-v0.1) jiang2024mixtral. However both of these models do not comply with the prompt format or refuse to answer.
@@
Phi refuses to answer in the required format in the majority of cases. In a test run of 544 * 3 = 1632 inferences, the model refused to comply with the format in 717 of them. In 110 cases the model outright refuses to answer with the phrase “As an AI language model, …”. Therefore we do not pursue Phi-mini further.
@@
Out of 27700 votes * 3 runs = 83310 inferences, Mixtral refuses to answer in 86 votes (we prompt multiple times to account for inconsistency of the output. In these cases the model does not comply even when repeatedly prompted). Additionally, it does not comply with the answer restrictions in 275 cases, giving different answers e.g. “Undecided”. Disregarding such cases, the model achieves a weighted F1-Score of 0.698 when prompted with wikipedia prompts and reasoning, which is comparable to the performance of Llama3-8B with wikipedia prompts and no reasoning. Due to the increased computational costs, because of the larger model size, repeated prompting, as well as the worse performance, partly due to the non-compliance of the answer format, we do not conduct further experiments with the model.
Appendix G Prompting Examples
@@
We show our different approaches to prompting in Figure 7. We either construct the persona with a summary of the Wikipedia article of the politician generated with Llama3-70B, or we construct the persona with the various attributes, consisting of the full name, the gender, age, birthplace and the country, national party and the group of the politician. Then we give the proposal information, either with the real speeches of the European Parliament or the counterfactual ones created with Llama3-70B. Then we instruct the model to respond in JSON-Format and to either directly respond with the vote or to first generate a reasoning chain before casting a vote.
@@
Examples of responses are presented in Figure 8. Models frequently mention the national party and European group associated with their persona, and sometimes reference their nationality. To the best of our knowledge, the reasoning chains generally align with the final vote; however, we do not conduct a qualitative analysis.
@@
We show two real speeches of the European Parliament and the counterfactual speeches generated by the model in Figure 9. We use the prompt In the following you are given a speech arguing for, against or being neutral towards a proposal in the EU parliament with the name [Title]. First, indicate which stance the given speech takes. Then, generate a new speech that takes the opposite stance (or stays neutral) and only use arguments and information that appear in the given speech. to determine the stance the LLM has towards the speech and to generate the counterfactual speech. To verify whether the generated counterfactual speeches accurately reflect the opposite stance, three voluntary annotators rate a randomly chosen speech for each of the 47 proposals. We give the voluntary annotators the same first sentence as instruction to determine the stance of the speech. Based on a majority vote, 43 out of 47 speeches indeed align with the opposite stance relative to the original speeches, with a high level of agreement among annotators (Fleiss’ Kappa score of $0.800$).
@@
Appendix H Compute Infrastructure and Time
We utilize a dual-server setup, each equipped with two NVIDIA H100 GPUs (80 GB VRAM) and two AMD EPYC 9124 16-Core processors. Due to our large input length, generating reasoning chains significantly increases the computational time required for each simulation. On average, a full simulation without reasoning takes approximately 3.5 hours for Llama3-8B and Qwen-7B on a single GPU, around 18 hours for Llama3-70B and 21 hours for Qwen-72B on two GPUs. However, when reasoning is included, the runtime increases to roughly 17 hours for Llama3-8B and Qwen-7B on a single GPU and approximately 81 hours for Llama3-70B and 95 hours for Qwen-72B on two GPUs. We approximate another 100 dual-GPU hours for testing purposes and creating the counterfactual speeches.

### v3

@@ Appendix E Counterfactual Debates
Appendix F Proposals and Committees
We show the full list of names of the proposals and the committees that initiated them in Table 9. We selected these proposals because they had a significant societal impact, evidenced by an official EU press release focused on that proposal decision and by the fact that they were the sole focus of one or more debates in the European Parliament. We analyzed the group line for each proposal and compared it with the predicted group line of our best performing approach Llama3-70B using reasoning and attribute prompts. We show the comparison in Table 5. We see that the model never predicts the ABSTENTION group line. The Renew group was predicted best, followed by S&D, Greens/EFA and EPP. For the far-left group GUE/NGL, the model predicts more FOR group lines on topics that include Ukraine and policies that introduce more digital processes, which involve more data collection of citizens (e.g., European Health Data Space or Instant Payments in Euro). For the ECR group, the model predicts more agreement with proposals that aim to preserve nature and reduce emissions (e.g., Nature Restoration, Methane emissions reduction in the energy sector). The ID group has the group line of ABSTENTION most often, which the model did not predict at all. Further false predictions are again related to ecological proposals; however, FOR this group, the model predicts more AGAINST group lines than FOR group lines (e.g., Ozone Depleting Substances, Empowering consumers for the green transition). The vote breakdown indicates that our current models struggle to predict when politicians from centrist parties vote against their group’s established position. Conversely, for groups on the political fringes, the models sometimes fail to predict the group line itself.
@@ Appendix F Proposals and Committees
We analyzed the change in voting behavior across different policy categories when using counterfactual speeches as well. To explore this, we categorize proposals based on the committee that initiated them. Figure 6 reveals that individuals from left-leaning (S&D, Greens/EFA) and even far-left groups (GUE/NGL) consistently maintain their support for proposals on certain topics, such as those from the Employment and Social Affairs Committee and the Women’s Rights and Gender Equality Committee. For these proposals, most original speeches supported them ($83\%$ and $72\%$, respectively), so the majority of counterfactual debates advise against them. Despite this, the votes of individuals affiliated with left-wing groups remain unchanged. In contrast, center-right (Renew, EPP) and far-right groups (ECR, ID) are more likely to alter their votes on proposals from these committees.
@@
Appendix G Influence of different attributes
@@
G.1 Attribute Matching
We make use of regular expressions to determine which attributes appear in the reasoning chains of each MEP and to remove specific words from speeches. In the following, we list the substrings used to search for the values of each attribute in the reasoning chains. For each MEP, we only search for their actual attribute values (i.e., we search for Ireland and Irish if a MEP represents Ireland in the EU parliament). For the attribute age, we computed the age of each MEP at the time of voting. We used the exact wording for birthplace , which comes from Wikidata [^6] https://www.wikidata.org/. We also used the exact wording for the politician’s name and national party from the European Parliament Website [^7] https://www.europarl.europa.eu/meps/en/full-list/all.
@@ G.1 Attribute Matching
G.2 Additional Attribute Analysis
@@
Table 6 shows the gender in the reasonings. We disaggregate the mentions of gender identifiers by the gender of the MEPs (e.g., gender_fm indicates the percentage of reasonings by female (f) simulated MEPs that mention a male (m) identity word). We find that the percentages remain roughly consistent across all prompting strategies and that female identity words are mentioned much more often than male identity words. Further, we observe that the reasoning chains contain male and female identity words at a similar rate, regardless of the gender in the prompt. We conclude that in the reasoning chains, gender identity terms are likely not used to refer to the gender of the persona (the simulated "own" gender), but rather to discuss a proposal related to gender (e.g., "Combating violence against women and domestic violence") or other proposals initiated by the Committee of Women’s Rights and Gender Equality.
@@
Table 7 presents the number of politicians for whom the correct group is mentioned at least once in the reasoning when only their names are provided as personas. The pre-training data appears to contain information on a significant portion of politicians; however, the distribution is biased towards center parties. Although the pre-training data likely contains more information than what is provided in the prompt, we still observe a significant change in behavior when including all attributes in the prompt instead of just the name. This suggests that using only the name is not always sufficient to guide the model toward a known public figure.
Appendix H Other models
@@
To prevent knowledge leakage we want to restrict used models to pretraining data cutoff before the date of our earliest vote in our dataset (16. January 2024) to the best of our abilitites and have comparable performance to LLama-3 and Qwen2.5. We identify two other models that fit this criteria: Phi-mini (Phi-3.5-mini-instruct Abdin et al. (2024) and (Mixtral-8x7B-Instruct-v0.1) Jiang et al. (2024a). However, neither of these models complies with the prompt format nor refuses to answer.
@@
Phi refuses to answer in the required format in the majority of cases. In a test run of 544 * 3 = 1632 inferences, the model refused to comply with the format in 717 of them. In 110 cases, the model outright refuses to answer with the phrase “As an AI language model, …”. Therefore, we do not pursue Phi-mini further.
@@
Out of 27700 votes * 3 runs = 83310 inferences, Mixtral refuses to answer in 86 votes (we prompted multiple times to account for the inconsistency of the output. In these cases, the model does not comply, even when repeatedly prompted). Additionally, it does not comply with the answer restrictions in 275 cases, giving different answers, e.g., “Undecided”. Disregarding such cases, the model achieves a weighted F1-Score of 0.698 when prompted with wikipedia prompts and reasoning, which is comparable to the performance of Llama3-8B with wikipedia prompts and no reasoning. Due to the increased computational costs because of the larger model size, repeated prompting, and worse performance, partly due to the non-compliance with the answer format, we do not conduct further experiments with the model.
Appendix I Prompting Examples
@@
We show our different approaches to prompting in Figure 9. We either construct the persona using a summary of the Wikipedia article of the politician generated with Llama3-70B, or we construct the persona with various attributes, consisting of the full name, gender, age, birthplace, country, national party, and group of the politician. Then we provide the proposal information, either with the real speeches of the European Parliament or the counterfactual ones created with Llama3-70B. Then we instruct the model to respond in JSON-Format and either to respond directly with the vote or to first generate a reasoning chain before casting a vote.
@@
Examples of responses are presented in Figure 7. Models frequently mention the national party and European group associated with their persona, and sometimes reference their nationality.
@@
We show two real speeches from the European Parliament and the counterfactual speeches generated by the model in Figure 8. We use the prompt In the following you are given a speech arguing for, against or being neutral towards a proposal in the EU parliament with the name [Title]. First, indicate which stance the given speech takes. Then, generate a new speech that takes the opposite stance (or stays neutral) and only use arguments and information that appear in the given speech. to determine the stance the LLM has towards the speech and to generate the counterfactual speech. To verify whether the generated counterfactual speeches accurately reflect the opposite stance, three voluntary annotators rate a randomly chosen speech for each of the 47 proposals. We give the voluntary annotators the same first sentence as instruction to determine the stance of the speech. Based on a majority vote, 43 out of 47 speeches indeed align with the opposite stance relative to the original speeches, with a high level of agreement among annotators (Fleiss’ Kappa score of $0.800$).
@@
Appendix J Compute Infrastructure and Time
We utilize a dual-server setup, each equipped with two NVIDIA H100 GPUs (80 GB VRAM) and two AMD EPYC 9124 16-Core processors. On average, a full simulation without reasoning takes approximately 3.5 hours for Llama3-8B and Qwen-7B on a single GPUs, around 18 hours for Llama3-70B and 21 hours for Qwen-72B on two GPUs. However, when reasoning is included, the runtime increases to roughly 17 hours for Llama3-8B and Qwen-7B on a single GPU and approximately 81 hours for Llama3-70B and 95 hours for Qwen-72B on two GPUs. We approximate another 100 dual-GPU hours for creating the counterfactual speeches.
Figure 9: Prompting example. This shows the different approaches we take to prompt the model. We give a persona description, which is either a Wikipedia Summary or an attribute prompt, proposal information in the form of debates (only real speeches are depicted here) and instructions on how to format the response.